OpenStudy (anonymous):

The square root of -8i is?

OpenStudy (amriju):

8^.5(1+i)

OpenStudy (anonymous):

sorry it's wrong

OpenStudy (amriju):

yeah 2(1+i)

OpenStudy (anonymous):

The answer is \[\pm 2(1-i)\] please explain ,how?

OpenStudy (amriju):

1-i?..r u sure?

OpenStudy (anonymous):

yes i am damn sure

OpenStudy (amriju):

pellet..its -8i...sorry..i did for 8i..yeah its correct

OpenStudy (anonymous):

How> explain the steps?

OpenStudy (amriju):

I'll telll u the easiest process...take a+ib to be the root..square it then equate the real and imaginary parts with -8i...find a and b

OpenStudy (anonymous):

I tried the same method....but i didn't get it. please write the steps you have taken.

OpenStudy (amriju):

(a+ib)^2=-8i a^2+2iab-b^2=-8i.. now real part is 0..so a^2=b^2...meaning a=+/- b equating imaginary parts 2ab=-8 now if a=b, then 2a^2=-8..thats not possible( and b are obviously real) so a=-b so -2a^2=-8 a=2 b=-2 so 2(1-i)

OpenStudy (amriju):

or else try using eulers method..and de moivre's theorem...thats the fastest process

OpenStudy (anonymous):

Thanks @amriju

OpenStudy (amriju):

w.c. :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Euler is much easy.....

OpenStudy (amriju):

yup..and its fastest...

OpenStudy (anonymous):

ok one more please....

OpenStudy (amriju):

okk..n.p

OpenStudy (amriju):

i'll try

OpenStudy (anonymous):

square root of 3+4i?? before one i tried from euler formula i got the answer but this won't come from that.

OpenStudy (amriju):

why don't you try a+ib..? ok..if u want to try euler 5(3/5+4/5i) u know how I did it..dont u? 5 (cos 53 +i sin 53) 5e^(i*53pi/180) sq. root it sq.root(5)*e^(i*53pi/360) now convert it into cos x + isin x form..and solve

OpenStudy (anonymous):

ok i got it...thanks ok can you give me procedure of a+ib .... i am confused somewhat in it.

OpenStudy (amriju):

the first step is to assume a root of the form a+ib square it to get a^2+2iab -b^2 now look at the rhs...it has a real and a complex part, rite real and complex parts never mess with one another..so a^2-b^2=real part in the rhs 2iab=complex part of rhs.. two variables, two eqns..solve it for a and b

OpenStudy (amriju):

can you do it now?

OpenStudy (anonymous):

If i have some, i will ask... thanks btw